while training for an iron man competition, tony bikes for 48 miles and runs for 12 miles. if his biking speed if more than 8 times his running speed and it take 4 hours to complete the training, how long did he spend on his bike?
Answer by Boreal(15235)

running speed =x
bike speed =8x

time to run = 48/8x
time to run= 12/x
(48/8x) + (12/x) = 4 units miles/miles/hour = hours
Multiply by 8x, both sides
48 + 96 = 32x
144=32x
x=4.5 mph
8x=36 mph
Takes him 1 1/3 hours to bike 48 miles ANSWER
Takes 2 2/3 hours to run 12 miles.
Note, I let bike speed be 8 times the running speed, not "more than," which would change these numbers.